Solve by completing the square.
x^2 - 16x = 10
I tried using one of the solvers and it didn't seem to work for this problem.

Answer by edjones(8007)

x^2-16x=10
x^2-16x+64=10+64 take (-16/2=-8) and square it (64) and use it as the third number. (completing the square.
(x-8)^2=74
x-8=+-sqrt(74) Sqrt of both sides
x=8+-sqrt(74)
x=8+sqrt(74), x=8-sqrt(74)
.
Ed